OpenStudy (campayne):

A restaurant charges $8 for a large cheese pizza plus $1.50 per topping. Another restaurant charges $11 for a large cheese pizza plus $0.75 per topping. To find out how many toppings it would take for the pizzas to cost the same at both restaurants, which equation would you set up? A) 8x + 1.5 = 0.75x + 11 B) 8 + 0.75x = 11 + 1.5x C) 8 + 1.5x = 11 + 0.75 x D) 8x + 1.5 = 11x + 0.75

OpenStudy (imnotarobot5):

So the first equation will have the base cost at 8 dollars and the variable cost at 1.50

OpenStudy (imnotarobot5):

so do you know what the first equation would be?

OpenStudy (campayne):

umm. sounds like C

OpenStudy (imnotarobot5):

Well that wont give you the answer quite yet

OpenStudy (imnotarobot5):

Your equation for the first pizza is 8 + 1.50x

OpenStudy (imnotarobot5):

Does that make sense?

OpenStudy (campayne):

indeed

OpenStudy (imnotarobot5):

Alright so the equation for the second pizza is 11 + .75x

OpenStudy (imnotarobot5):

In order to find out how many toppoings it'd take for both to be equak, you set the two equations equal to eachother

OpenStudy (imnotarobot5):

8 + 1.50x = 11 + .75x

OpenStudy (imnotarobot5):

Then solve accordingly

OpenStudy (campayne):

oh boy.. well here we go

OpenStudy (campayne):

I keep getting C

OpenStudy (imnotarobot5):

Haha you're right

OpenStudy (imnotarobot5):

Sorry, could have gotten you that earlier, I just thought you had to simplify

OpenStudy (campayne):

ahh okay well thanx :)

OpenStudy (campayne):

Now for part 2: For the same pizza restaurants referred to in the last question, how many toppings does a pizza have that costs the same at both restaurants? 3 4 8 10

OpenStudy (imnotarobot5):

1.50x - .75x = 11 - 8 .75x = 3

OpenStudy (imnotarobot5):

3/.75 = 4

OpenStudy (campayne):

yup i got it

OpenStudy (imnotarobot5):

Alright sweet